OpenStudy (gabby_taylor1224):

HELP! The equation of line AB is y+3x-2. Write an equation of a line parallel to line AB in slope-intercept form that contains point (-2,4)

OpenStudy (gabby_taylor1224):

@Etrainx do you know how to do this?

OpenStudy (gabby_taylor1224):

@jojobuck

OpenStudy (gabby_taylor1224):

@Vocaloid

Vocaloid (vocaloid):

is the equation y = 3x-2 or y+3x=-2 or something else?

OpenStudy (anonymous):

i'm pretty sure it's y = 3x-2

OpenStudy (anonymous):

At least that would make the most sense

OpenStudy (gabby_taylor1224):

@Vocaloid im not sure this was all the question said

Vocaloid (vocaloid):

ok, assuming our equation is y = 3x-2 and our point is (-2,4) parallel means same slope, so our slope is 3 then we use the point slope equation y-y1 = m(x-x1) where m is the slope and (x1,y1) is our given point